# See pages 1 of figures/front.pdf

### v1

@@ 2.1 Real-Time Analysis
Real-time analysis, as the name suggests, is the processing of data in “real time”. However, its definition varies widely between disciplines and even within these disciplines and their specific use cases. In general, these systems are of interest in scenarios where making a decision is constrained in some way by time. The manner of this constrain varies between the different scenarios. For example, on the one hand, a system developed for the control of self-driving cars, in order to ensure the safety of the passengers, may need to have a reaction time below a specific threshold, for example lower than a peak-physique human driver. This response time is known as latency and we would say that this system for autonomous vehicles is latency-constrained. This latency constraint is known as a hard real-time constraint. As defined in [36], “A real-time constraint is called hard, if not meeting that constraint could result in a catastrophe”—car accidents that may even be fatal.
On the other hand, for example for multimedia streaming applications, timely processing is preferable, but delayed processing does not cause system failure or result in catastrophic consequences. However, failing to meet time expectations can lead to reduced output quality, such as the freezing of the frames of the video etc. This type of real-time constraint is known as soft.
@@
Many of these data processing systems are indeed latency-constrained. Hardware triggers, specifically, have to decide whether to keep or discard the incoming data every few tens of nanoseconds: otherwise, the data are permanently lost. This hard time limit, is what ultimately constrains how much processing you can pack into the trigger. By contrast, other architectures prioritize throughput—the amount of data processed per unit time. In software triggers, for example, the incoming data from the detectors have to be processed by the system in order to make a decision about whether they should be saved or not, and, as before, a failure of this decision could result in the permanent loss of the data. Assuming the data come in streams of data packets, known in HEP as events, a constraint on throughput is equivalent to a constraint on the average processing time of each data packet. As long as the average processing time is below a specific threshold, the processing time of a single data packet can exceed that threshold. These systems are known as throughput-constrained.
@@
    • Threat Detection: Survellaince of network traffic in order to detect potential secutiry threats [59].
@@ 2.2 Real-Time Analysis in High-Energy Physics
Specifically for LHCb, the experiment will operate at its current configuration until the end of Run 4 (2033) reaching a maximum instantaneous luminosity of $L=2\times 10^{33}~\text{cm}^{-2}\,\text{s}^{-1}$ [64]. By the end of of Run 4 it will have recorded 50 fb-1 of high-energy pppp collision data. In contrast, after the upgrade during Long Shutdown 4 (LS4), the detector will operate at $L=1\times 10^{34}~\text{cm}^{-2}\,\text{s}^{-1}$, corresponding to $\langle\mu\rangle\approx$ 28–42 interactions per bunch crossing, compared to the current $\langle\mu\rangle\approx 5$. By the end of the HL-LHC operation, the detector will have recorded at least 300 fb-1. A summary of the integrated luminosities is shown in Fig. 2.3. Furthermore, Fig. 2.4 illustrates how data bandwidth has evolved over time across past and upcoming experiments. This demonstrates the scale of the challenge currently facing the LHC community.
@@
In order for the HL-LHC project to be successfully completed, a massive effort is underway. The detectors and all the infrastructure at the LHC—including magnets, cryogenics, vacuum systems, and beam instrumentation—has be to be modernized.
@@
Firstly, hardware triggers are fast and simple, but operate on coarse detector data and hence on a rudimentary representation of the collision events. On the other hand, RTA allows a more sophisticated and physics-rich event selection using information from various subdetectors. In this way, interesting events, that would otherwise have been discarded by the coarse filtering of the hardware trigger, can still be kept. For example, rare processes traditionally studied by the LHCb experiment, such as Charge-Conjugation Parity (CP) violation and heavy flavor decays, may not trigger efficiently on hardware but can be identified well using full tracking information in real time.
@@ 2.3 RTA using ML on Heterogeneous Architectures
Furthermore, ML applications “on the edge”, i.e., ultra-low-latency and on-detector/sensor, are growing day by day. Examples include $\tau$ [24], $b$-quark [78], and electron [79] identification, anomaly detection [80, 81, 82], data compression [83], and continual learning [84] in the CMS trigger. Other examples include calorimeter peak finding [85] in the ATLAS trigger.
@@ 3.1 Accelerator Physics
In experimental particle physics, another frequently used spatial coordinate is the pseudorapidity $\eta$ . It describes the angle between a particle’s momentum $\mathbf{p}$ and the positive direction of the beam axis—identified with the $z$-direction. This angle is refereed to as the polar angle $\theta$, as shown in Fig. 3.2.
@@ 4.1 Machine Learning
Machine learning is the field of how machines—specifically computers—can “learn”. Although “learn” is perhaps a generous term, it refers to how computers manage to do specific tasks without being explicitly programmed to do them. Unlike classical algorithms, which follow hand-crafted rules defined by developers, ML algorithms, and by consequence ML models, are data-driven: By an iterative process of providing data to the ML model, the model is trained and progressively learns to perform a task solely based on the data it has be given. At the end of this process, without the need for the developer to describe the logic of the algorithm itself, the model can carry out the task effectively without the developer needing to explicitly define how it should be done.
@@
Classical, or probabilistic, ML have been in use long before the term ML came into existence. These algorithms are statistical models who are trying to capture relationships between various variables. Arguably, the most famous example is linear regression, originally developed by Isaac Newton for his work on the equinoxes around 1700 [131], and later formalized by Legendre and Gauss in the early 19th century [132].
@@
However, what happens if we pass each photo through a color filter, changing the color of the pixels? In this case, the algorithm breaks down completely. However, to a human eye, the classification task remains identically easy. So, how do we capture the “seaness” of the sea and the “grassness” of the grass? This is exactly where things get difficult. It is not obvious how to design a feature exactly in order to capture, for example, the texture of the grass in terms of pixel values. This is where representation learning, also known as feature learning, comes in. It is a set of techniques that allows a system to automatically discover the representation needed for a specific problem, completely bypassing the need for hand-designing. And as it turns out, learned representations often result in much better performance that hand-designed ones [76].
@@
Another reason why deep learning has only recently gained such traction is the availability of data in the era of “big data”. ML algorithms are data-driven and in fact need a large amount of data in order to be able to be trained and to generalize well on unseen data. With the increasing digitization of society, data became abundant. Furthermore, it was possible to gather all this records and curate them into large datasets appropriate for training ML models.
@@
Other learning paradigms exists such as semi-supervised learning and reinforcement learning. The former is when some examples in the dataset include supervision targets while others do not, while the latter is when the learning algorithm interacts with an environment, so there is a feedback loop between the learning system and its actions.
@@
$=\frac{1}{m}\sum_{i=1}^{m}(\hat{\mathbf{y}}-\mathbf{y})_{i}^{2}$ (4.4)
where $\mathbf{y}$ are the regression targets, and $m$ is the size of the set over which we are doing this evaluation. Furthermore, because we want to do a fair evaluation, we want to evaluate our model on examples it has never seen before. This can be achieved my splitting the dataset into a test and a train set. During the learning procedure the algorithm only has access to the training set, and after the end, the model is evaluated solely on the test set.
@@
This, given that $\epsilon_{i}=y_{i}-\mathbf{w}^{\top}\mathbf{x}_{i}$ from Eq. 4.12, implies that
@@ 4.2 Deep Learning
The hyperbolic tangent is defined as $\tanh(x)=(e^{x}-e^{-x})/(e^{x}+e^{-x})$ so the function takes values between 1-1 and 1. The function is zero-centered which can help with convergence compared to the sigmoid. Nonetheless, is still suffers from vanishing gradients for large inputs.
@@
The next question that arises is the following. Since we said our neural network is essentially a complex nested function of these combinations of nonlinear activations and affine transformations, as in Eq. (4.26), that means that the loss functions is going to have a similar structure. So, how do we know how to update the individual parameters of each layer of the neural network, in order to minimize this objective function?
@@
When we use a feedforward neural network that accepts an input $\mathbf{x}$ and produces an output $\mathbf{y}$, information flows “forward” through the network, as in from left to right in Fig. 4.4. The input vector $\mathbf{x}$ provides the initial information that propagates, layer by layer and finally results in $\mathbf{y}$. This vector $\mathbf{y}$ is a function of all the weights and biases of all the layers of the neural network, denoted collectively as $\boldsymbol{\theta}$. This process is known as forward propagation. A scalar cost function $J(\boldsymbol{\theta})$ can then be formed using the output $\mathbf{y}$.
@@ 4.3 Convolutional Neural Networks
This operation is performed for a number $k$ of filters, in order to extract various features in the image, and each filter’s parameters are completely independent. The output for each filter is different, and hence the operation of this convolutional layer results in a collection of $k$ feature maps. This collection can be thought of a higher-dimensional tensor and is called a volume. For color images, the input is actually also a volume, since the image is usually represented by three channels: R (red), G (green), and B (blue), where each channel is a monochrome picture.
@@ 4.4 Graph Neural Networks
Networks are ubiquitous—and so are graphs. In many real-world scenarios, its beneficial to think of data points not in isolation but as part of a web of complex connections: people connected through social interactions, proteins by biochemical interactions, or web pages by hyperlinks. Capturing and using this connectivity is crucial for understanding the underlying relationships and dynamics [159, 160, 161, 162].
@@
    The embeddings of neighboring nodes are aggregated using a permutation invariant function. This is justified because a permutation of the graph nodes should not give a different result. Examples of these aggregating functions include the max, sum or average functions. This process is refereed to as the aggregation of the messages received from the immediate neighbors.
@@ 4.5 Quantization
  • QAT: The range of computations are computed during training. “Fake quantize” operators simulate the effects of quantization during training, enabling the model to adjust and become robust to the errors introduced by the quantization process.
@@ 5.1 Parallelism
Ideally, doubling the number of processors would result in the halving of the runtime. However, in practice, very few algorithms achieve optimal speedup. The maximum potential speedup is given by Amdhal’s law [180]. A task executed on a multicore system can be categorized into two parts: a part that does not benefit from the usage of multiple cores, and a part that does benefit. Assuming that the latter is a fraction $\tau$ of the task, and that it benefits from an acceleration by a factor $s$ compared to single core execution then, the maximum speedup is given by:
@@
The relationship is illustrated in Fig. 5.1. Interestingly, this law reveals that increasing the number of processing yields diminishing returns past a certain point. In addition, it demonstrates that the enhancements of the code have to be focused on both the parallelizable and non-parallelizable components. Of course, the simplistic view of computation needed for the derivation of Amdahl’s law neglects various aspects of inter-process communication, synchronization and memory access overheads. A more complete assessment is given by Gustafson’s law [181].
@@
In order to address the problem of power consumption, manufacturers turned to producing power efficient processors that have multiple cores. Each core is independent and can access the same memory concurrently. This design principle brought multi-core processors to the mainstream. By early 2010s, computers by default had mutliple cores, while servers had more than ten core processors. By contrast, in early 2020s, some processors had over one hundred cores [179]. Moore’s law [185], that predicts that the number of transistors in an integrated circuit will double every roughly two years, can be extrapolated to the doubling of the number of cores per processor.
The operating system of the CPU ensures that the different tasks are performed concurrently using the resources of the processor by distributing them across the free cores. However, in order to unlock the full capacity of the processing unit, the code itself has to be designed in a way that leverages the new computational capabilites of multicore architectures [179].
@@ 5.2 From Video Games to the GPU Architecture
Early 3D graphics emerged in the 1990s in arcades and consoles and GPUs started integrating 3D functions. The term GPU was coined by Sony in reference to their 32-bit Sony GPU used in the PlayStation 1 video game console, released in 1994 [193]. Nvidia and ATI started creating consumer graphics accelerators, leading to the release of GeForce 256. This GPU was marketed as the world’s first GPU capable of performing advanced graphics rendering. These capabilities included tasks such as rasterization, where an image described in a vector graphics format is translated into a array of pixels that best represents this vector description in the available screen granularity. Shading, another essential task for a graphics processor, is the process through which a GPU calculates the appropriate levels of light and color, in order to render a 3D scene more realistically. The first GPU capable of shading was the GeForce 3, used in the Xbox console, competing with the chip used in PlayStation 2.
@@
Nvidia introduced the Compute Unified Device Architecture (CUDA) in 2006, sparking what is now known as General-Purpose Graphics Processing Unit (GPGPU) computing [145]. This marked a revolution in computing: previously, GPUs were dedicated chips designed to accelerate 3D rendering tasks for gaming and graphics applications. With CUDA, GPUs became programmable parallel processors equipped with hundreds of processing elements, enabling them to perform a broad range of tasks traditionally tackled using CPUs. This can include scientific computing (simulations, climate etc.), financial modeling, signal processing, machine learning and deep learning. For the first time, Nvidia provided a dedicated programming model and language for its GPUs, enabling developers to write general-purpose code that could run directly on the GPU—something that was previously not possible with such flexibility and ease.
@@
CUDA is a proprietary language, which led to the need for a standardized parallel programming language that could be used across GPUs from different manufacturers. In response, OpenCL [194, 195] was defined by Khronos Group as an open standard. It allows the development of code compatible with both GPU and CPU. This emphasis on portability—the ability to write a single kernel that can run across heterogenous platforms—made OpenCL the second most popular HPC tool at the time [196].
@@
In the 2010s, GPUs were used in consoles such as the PlayStation 4 and the Xbox One [197], and on automotive systems, after Nvidia partnered with Audi to power car dashboard displays [198]. Nvidia architectures developed further, increasing the number of CUDA cores and further adding the new technology of the so-called tensor cores [199]. Tensor cores were designed to bring better performance to deep learning operations. Real-time ray tracing—simulation of reflections, shadows, depth of field etc.—debuted with Nvidia RTX 20 series in 2018 [200].
@@
In 2020s, after the deep learning explosion we described in Chapter 4, GPUs are heavily used in the training an inference of large language models, such as the ChatGPT [150] chatbot by OpenAI. This surge in interest of dedicated hardware, infrastructure and electricity to support these heavy models has created a booming artificial intelligence ecosystem. It is further fueling a re-evaluation of our electricity needs, infrastructure organization, and the direction of hardware development, while also raising questions about the feasibility of continued scaling.

### v2

@@ 2.1 Real-Time Analysis
Real-time analysis, as the name suggests, is the processing of data in “real time”. However, its definition varies widely between disciplines and even within these disciplines and their specific use cases. In general, these systems are of interest in scenarios where making a decision is constrained in some way by time. The manner of this constraint varies between the different scenarios. For example, on the one hand, a system developed for the control of self-driving cars, in order to ensure the safety of the passengers, may need to have a reaction time below a specific threshold, for example lower than a human driver. This response time is known as latency and we would say that this system for autonomous vehicles is latency-constrained. This latency constraint is known as a hard real-time constraint. As defined in [36], “A real-time constraint is called hard, if not meeting that constraint could result in a catastrophe”—car accidents that may even be fatal.
On the other hand, for example for multimedia streaming applications, timely processing is preferable, but delayed processing does not cause system failure or result in catastrophic consequences. However, failing to meet time expectations can lead to reduced output quality, such as the freezing of the frames of the video, etc. This type of real-time constraint is known as soft.
@@
Many of these data processing systems are indeed latency-constrained. Hardware triggers, specifically, have to decide whether to keep or discard the incoming data every few tens of nanoseconds: otherwise, the data are permanently lost. This hard time limit is what ultimately constrains how much processing you can pack into the trigger. By contrast, other architectures prioritize throughput—the amount of data processed per unit time. In software triggers, for example, the incoming data from the detectors have to be processed by the system in order to make a decision about whether they should be saved or not, and, as before, a failure of this decision could result in the permanent loss of the data. Assuming the data come in streams of data packets, known in HEP as events, a constraint on throughput is equivalent to a constraint on the average processing time of each data packet. As long as the average processing time is below a specific threshold, the processing time of a single data packet can exceed that threshold. These systems are known as throughput-constrained.
@@
    • Threat Detection: Surveillance of network traffic in order to detect potential security threats [59].
@@ 2.2 Real-Time Analysis in High-Energy Physics
Specifically for LHCb, the experiment will operate at its current configuration until the end of Run 4 (2033) reaching a maximum instantaneous luminosity of $L=2\times 10^{33}~\text{cm}^{-2}\,\text{s}^{-1}$ [64]. By the end of Run 4 it will have recorded 50 fb-1 of high-energy pppp collision data. In contrast, after the upgrade during Long Shutdown 4 (LS4), the detector will operate at $L=1\times 10^{34}~\text{cm}^{-2}\,\text{s}^{-1}$, corresponding to $\langle\mu\rangle\approx$ 28–42 interactions per bunch crossing, compared to the current $\langle\mu\rangle\approx 5$. By the end of the HL-LHC operation, the detector will have recorded at least 300 fb-1. A summary of the integrated luminosities is shown in Fig. 2.3. Furthermore, Fig. 2.4 illustrates how data bandwidth has evolved over time across past and upcoming experiments. This demonstrates the scale of the challenge currently facing the LHC community.
@@
In order for the HL-LHC project to be successfully completed, a massive effort is underway. The detectors and all the infrastructure at the LHC—including magnets, cryogenics, vacuum systems, and beam instrumentation—has to be modernized.
@@
Firstly, hardware triggers are fast and simple, but operate on coarse detector data and hence on a rudimentary representation of the collision events. On the other hand, RTA allows a more sophisticated and physics-rich event selection using information from various subdetectors. In this way, interesting events that would otherwise have been discarded by the coarse filtering of the hardware trigger, can still be kept. For example, rare processes traditionally studied by the LHCb experiment, such as Charge-Conjugation Parity (CP) violation and heavy flavor decays, may not trigger efficiently on hardware but can be identified well using full tracking information in real time.
@@ 2.3 RTA using ML on Heterogeneous Architectures
Furthermore, ML applications “on the edge”, i.e., ultra-low latency and on-detector/sensor, are growing day by day. Examples include $\tau$ [24], $b$-quark [78], and electron [79] identification, anomaly detection [80, 81, 82], data compression [83], and continual learning [84] in the CMS trigger. Other examples include calorimeter peak finding [85] in the ATLAS trigger.
@@ 3.1 Accelerator Physics
In experimental particle physics, another frequently used spatial coordinate is the pseudorapidity $\eta$ . It describes the angle between a particle’s momentum $\mathbf{p}$ and the positive direction of the beam axis—identified with the $z$-direction. This angle is referred to as the polar angle $\theta$, as shown in Fig. 3.2.
@@ 4.1 Machine Learning
Machine learning is the field of how machines—specifically computers—can “learn”. Although “learn” is perhaps a generous term, it refers to how computers manage to do specific tasks without being explicitly programmed to do them. Unlike classical algorithms, which follow hand-crafted rules defined by developers, ML algorithms, and by consequence ML models, are data-driven: By an iterative process of providing data to the ML model, the model is trained and progressively learns to perform a task solely based on the data it has been given. At the end of this process, without the need for the developer to describe the logic of the algorithm itself, the model can carry out the task effectively without the developer needing to explicitly define how it should be done.
@@
Classical, or probabilistic, ML has been in use long before the term ML came into existence. These algorithms are statistical models that try to capture relationships between various variables. Arguably, the most famous example is linear regression, originally developed by Isaac Newton for his work on the equinoxes around 1700 [131], and later formalized by Legendre and Gauss in the early 19th century [132].
@@
However, what happens if we pass each photo through a color filter, changing the color of the pixels? In this case, the algorithm breaks down completely. However, to a human eye, the classification task remains identically easy. So, how do we capture the “seaness” of the sea and the “grassness” of the grass? This is exactly where things get difficult. It is not obvious how to design a feature exactly in order to capture, for example, the texture of the grass in terms of pixel values. This is where representation learning, also known as feature learning, comes in. It is a set of techniques that allows a system to automatically discover the representation needed for a specific problem, completely bypassing the need for hand-designing. And as it turns out, learned representations often result in much better performance than hand-designed ones [76].
@@
Another reason why deep learning has only recently gained such traction is the availability of data in the era of “big data”. ML algorithms are data-driven and in fact need a large amount of data in order to be able to be trained and to generalize well on unseen data. With the increasing digitization of society, data became abundant. Furthermore, it was possible to gather all these records and curate them into large datasets appropriate for training ML models.
@@
Other learning paradigms exist such as semi-supervised learning and reinforcement learning. The former is when some examples in the dataset include supervision targets while others do not, while the latter is when the learning algorithm interacts with an environment, so there is a feedback loop between the learning system and its actions.
@@
$=\frac{1}{m}\sum_{i=1}^{m}(\hat{\mathbf{y}}-\mathbf{y})_{i}^{2}\,$, (4.4)
where $\mathbf{y}$ are the regression targets, and $m$ is the size of the set over which we are doing this evaluation. Furthermore, because we want to do a fair evaluation, we want to evaluate our model on examples it has never seen before. This can be achieved by splitting the dataset into a test and a train set. During the learning procedure the algorithm only has access to the training set, and after the end, the model is evaluated solely on the test set.
@@
This, given that $\epsilon_{i}=y_{i}-\mathbf{w}^{\top}\mathbf{x}_{i}$ from Eq. (4.12), implies that
@@ 4.2 Deep Learning
The hyperbolic tangent is defined as $\tanh(x)=(e^{x}-e^{-x})/(e^{x}+e^{-x})$ so the function takes values between 1-1 and 1. The function is zero-centered which can help with convergence compared to the sigmoid. Nonetheless, it still suffers from vanishing gradients for large inputs.
@@
The next question that arises is the following. Since we said our neural network is essentially a complex nested function of these combinations of nonlinear activations and affine transformations, as in Eq. (4.26), that means that the loss function is going to have a similar structure. So, how do we know how to update the individual parameters of each layer of the neural network, in order to minimize this objective function?
@@
When we use a feedforward neural network that accepts an input $\mathbf{x}$ and produces an output $\mathbf{y}$, information flows “forward” through the network, as in, from left to right in Fig. 4.4. The input vector $\mathbf{x}$ provides the initial information that propagates, layer by layer and finally results in $\mathbf{y}$. This vector $\mathbf{y}$ is a function of all the weights and biases of all the layers of the neural network, denoted collectively as $\boldsymbol{\theta}$. This process is known as forward propagation. A scalar cost function $J(\boldsymbol{\theta})$ can then be formed using the output $\mathbf{y}$.
@@ 4.3 Convolutional Neural Networks
This operation is performed for a number $k$ of filters, in order to extract various features in the image, and each filter’s parameters are completely independent. The output for each filter is different, and hence the operation of this convolutional layer results in a collection of $k$ feature maps. This collection can be thought of as a higher-dimensional tensor and is called a volume. For color images, the input is actually also a volume, since the image is usually represented by three channels: R (red), G (green), and B (blue), where each channel is a monochrome picture.
@@ 4.4 Graph Neural Networks
Networks are ubiquitous—and so are graphs. In many real-world scenarios, it is beneficial to think of data points not in isolation but as part of a web of complex connections: people connected through social interactions, proteins by biochemical interactions, or web pages by hyperlinks. Capturing and using this connectivity is crucial for understanding the underlying relationships and dynamics [159, 160, 161, 162].
@@
    The embeddings of neighboring nodes are aggregated using a permutation invariant function. This is justified because a permutation of the graph nodes should not give a different result. Examples of these aggregating functions include the max, sum or average functions. This process is referred to as the aggregation of the messages received from the immediate neighbors.
@@ 4.5 Quantization
  • QAT: The ranges of computations are computed during training. “Fake quantize” operators simulate the effects of quantization during training, enabling the model to adjust and become robust to the errors introduced by the quantization process.
@@ 5.1 Parallelism
Ideally, doubling the number of processors would result in the halving of the runtime. However, in practice, very few algorithms achieve optimal speedup. The maximum potential speedup is given by Amdahl’s law [180]. A task executed on a multicore system can be categorized into two parts: a part that does not benefit from the usage of multiple cores, and a part that does benefit. Assuming that the latter is a fraction $\tau$ of the task, and that it benefits from an acceleration by a factor $s$ compared to single core execution then, the maximum speedup is given by:
@@
The relationship is illustrated in Fig. 5.1. Interestingly, this law reveals that increasing the number of processors yields diminishing returns past a certain point. In addition, it demonstrates that the enhancements of the code have to be focused on both the parallelizable and non-parallelizable components. Of course, the simplistic view of computation needed for the derivation of Amdahl’s law neglects various aspects of inter-process communication, synchronization and memory access overheads. A more complete assessment is given by Gustafson’s law [181].
@@
In order to address the problem of power consumption, manufacturers turned to producing power efficient processors that have multiple cores. Each core is independent and can access the same memory concurrently. This design principle brought multi-core processors to the mainstream. By early 2010s, computers by default had multiple cores, while servers had more than ten core processors. By contrast, in early 2020s, some processors had over one hundred cores [179]. Moore’s law [185], that predicts that the number of transistors in an integrated circuit will double every roughly two years, can be extrapolated to the doubling of the number of cores per processor.
The operating system of the CPU ensures that the different tasks are performed concurrently using the resources of the processor by distributing them across the free cores. However, in order to unlock the full capacity of the processing unit, the code itself has to be designed in a way that leverages the new computational capabilities of multicore architectures [179].
@@ 5.2 From Video Games to the GPU Architecture
Early 3D graphics emerged in the 1990s in arcades and consoles and GPUs started integrating 3D functions. The term GPU was coined by Sony in reference to their 32-bit Sony GPU used in the PlayStation 1 video game console, released in 1994 [193]. Nvidia and ATI started creating consumer graphics accelerators, leading to the release of GeForce 256. This GPU was marketed as the world’s first GPU capable of performing advanced graphics rendering. These capabilities included tasks such as rasterization, where an image described in a vector graphics format is translated into an array of pixels that best represents this vector description in the available screen granularity. Shading, another essential task for a graphics processor, is the process through which a GPU calculates the appropriate levels of light and color, in order to render a 3D scene more realistically. The first GPU capable of shading was the GeForce 3, used in the Xbox console, competing with the chip used in PlayStation 2.
@@
Nvidia introduced the Compute Unified Device Architecture (CUDA) in 2006, sparking what is now known as General-Purpose Graphics Processing Unit (GPGPU) computing [145]. This marked a revolution in computing: previously, GPUs were dedicated chips designed to accelerate 3D rendering tasks for gaming and graphics applications. With CUDA, GPUs became programmable parallel processors equipped with hundreds of processing elements, enabling them to perform a broad range of tasks traditionally tackled using CPUs. This can include scientific computing (simulations, climate, etc.), financial modeling, signal processing, machine learning and deep learning. For the first time, Nvidia provided a dedicated programming model and language for its GPUs, enabling developers to write general-purpose code that could run directly on the GPU—something that was previously not possible with such flexibility and ease.
@@
CUDA is a proprietary language, which led to the need for a standardized parallel programming language that could be used across GPUs from different manufacturers. In response, OpenCL [194, 195] was defined by Khronos Group as an open standard. It allows the development of code compatible with both GPU and CPU. This emphasis on portability—the ability to write a single kernel that can run across heterogeneous platforms—made OpenCL the second most popular HPC tool at the time [196].
@@
In the 2010s, GPUs were used in consoles such as the PlayStation 4 and the Xbox One [197], and on automotive systems, after Nvidia partnered with Audi to power car dashboard displays [198]. Nvidia architectures developed further, increasing the number of CUDA cores and further adding the new technology of the so-called tensor cores [199]. Tensor cores were designed to bring better performance to deep learning operations. Real-time ray tracing—simulation of reflections, shadows, depth of field, etc.—debuted with Nvidia RTX 20 series in 2018 [200].
@@
In 2020s, after the deep learning explosion we described in Chapter 4, GPUs are heavily used in the training and inference of large language models, such as the ChatGPT [150] chatbot by OpenAI. This surge in interest of dedicated hardware, infrastructure and electricity to support these heavy models has created a booming artificial intelligence ecosystem. It is further fueling a re-evaluation of our electricity needs, infrastructure organization, and the direction of hardware development, while also raising questions about the feasibility of continued scaling.